# Entropic signature of resonant thermal transport: Ordered form of heat conduction

### v3

@@ Abstract
Thermal transport in crystals is influenced by chemistry, boundaries, and nanostructure. The anharmonic phonon band structure extracted from molecular-dynamics simulations provides an illuminating view of both the type and extent of prevalence of wavelike mechanisms underlying the transport, yet falls short of elucidating the nature of thermal evolution for different phonon regimes. Here we present an analysis framework for the characterization of the entropic signature of the mechanisms induced by boundaries and nanostructure, using both equilibrium and nonequilibrium atomistic simulations. Specifically, we examine the effects of phonon confinement, Bragg scattering, and local resonances on the configurational phase space in room-temperature nanostructured silicon, and quantify how each modifies the rate of entropy production and thermal relaxation. We reveal that the presence of phonon local resonances spanning the full spectrum enables a highly ordered regime of heat conduction, where irreversible evolution and entropy maximization are severely hindered by extensive mode hybridizations caused by the resonances. This unique regime of transport paves the way for ultra-precise phonon control for a wide range of applications in condensed matter physics.
@@ I Introduction
The second law of thermodynamics steers the spontaneous evolution of isolated material systems from arbitrary initial conditions towards the maximum entropy state corresponding to equilibrium. In crystal solids, Debye and Peierls, among others, conjectured that the anharmonic nature of atomic interactions combined with the presence of naturally-occurring lattice defects form the key causes behind the tendency of systems to thermally relax [1]. Consistent with this view, nonlinear phonon-phonon and phonon-defect scattering events are interpreted in the phonon gas picture to not necessarily conserve the lattice quasi-momentum, thus justifying on microscopic grounds the diffusion of vibrational energy along macroscopic thermal gradients. While all atomic interactions and motions are overseen by the Hamiltonian that governs a given system, the specific path and rate of evolution towards thermodynamic equilibrium may be affected by configurational constraints that drive certain mechanistic dynamical behavior. Such inherent competition between dynamical mechanisms and the emergence of equilibrium were brought to light by pioneering computational experiments reported by Fermi and colleagues [2]. Lattice disorder and its impact on localization provides an example of a key mechanism that influences the evolution to equilibrium [3, 4]. More recently, a number of other mechanisms that may hinder thermodynamic behavior have been investigated [5], which include interaction with thermal reservoirs [6, 7], phonon confinement in low-dimensional systems [8, 9], and nonlinear interfacial effects [10].
@@
Nanostructuring offers yet another rich platform for the disruption of thermal evolution. When feature sizes are confined to the range of the phonon mean free path (MFP) at a given temperature, the intrinsic structural dynamics become a key factor in potentially influencing the path to equilibrium. In this context, a distinct mechanism that has been shown to cause transformative changes to the phonon band structure is atomic-scale resonance hybridizations [11]. The presence of sub-MFP resonating substructures generates localized phonon resonances that couple with the underlying heat-carrying phonons traveling in the host medium [12]. This phenomenon may in principle take place across the full phonon spectrum, thus bringing rise to a unique form of heat conduction that is based on resonant thermal transport [13]. In this regime, spectral energy density (SED) calculations obtained from equilibrium molecular-dynamics (MD) simulations have revealed distinctly appearing flat dispersion curves that couple with the phonon dispersion throughout the full frequency range of atomic motion [14]. Nanostructures exhibiting this behavior, referred to as nanophononic metamaterials (NPMs), have recently been experimentally demonstrated [15]. Another wave-based phonon mechanism that also fundamentally influences the phonon band structure is Bragg scattering, which comes into effect with the presence of periodic features smaller than the phonon MFP [16, 17, 18, 19].
@@
The impact of the effects mentioned above on the emergence of thermal equilibrium has not yet been studied from a thermodynamics perspective. Yet analysis techniques and experiments have been developed for the characterization of thermodynamic processes in terms of entropy evolution [20]--a perspective that opens opportunities for exploring the interplay between structural dynamics and statistical mechanics [21, 22]. For example, generalized notions of entropy beyond local equilibrium have been applied towards the development of nanoscale thermal transport models not limited to diffusion [23, 24]. In addition, the equilibrium configurational entropy has been extracted directly from MD simulations [25, 26]. Noticeable applications include the characterization of phase transitions [27], defects [28, 29], biochemical phenomena [30], and the identification of universal scaling laws of transport properties in liquids [31]. Another key recent advancement is the emergence of thermodynamic uncertainty relations (TURs) [32, 33] built on fluctuation-dissipation theorems [34] as this provides new tools to relate the fluctuations of dissipative fluxes to the production of entropy in nonequilibrium processes [35, 36, 37].
@@
In this article, we quantify--for the first time--the entropic signature of non-diffusive heat flow in the presence of dynamical constraints created by material nanostructuring, and we do so using MD simulations. The resulting framework bridges the phonon dispersion properties of any given class of nanostructures with the thermodynamic picture and offers a new route to distinguishing the coherent behavior of phonons. We consider a variety of crystalline nanoscale features that influence multiple phonon transport mechanisms at room temperature, including boundary scattering, confinement, Bragg scattering, and local resonances. Considering silicon as an example constituent material, for each system we first obtain the phonon band structure by calculating the spectral energy density of atomic motion obtained by MD simulation where anharmonic effects are accounted for. The dispersion relations highlight the influence of the type of nanostructure on the wavelike nature of phonons. Next we characterize the configurational entropy distribution at equilibrium and uncover non-homogeneous contributions influenced by the presence of the nanoscale features. We then combine nonequilibrium simulations and TUR to obtain a lower bound for the entropy production rate. This metric quantifies the impact of inherent dynamical phenomena on the production and exchange of entropy with the external environment, which allows us to characterize the strength of each configurational constraint in inhibiting the progression towards equilibrium. We find that NPMs are exceptional in this aspect owing to the unique ability of full-spectrum phonon-resonance hybridizations in hindering irreversible evolution. The existence of this kind of resonant phonon behavior, which manifests as a strongly fluctuating energy current, demonstrates that thermal energy currents do not necessarily correspond to entropy currents at the nanoscale. This brings to light a remarkable observation of reversibility in phonon transport in the presence of dynamical constraints.
@@
For purposes of comparison, we consider the following geometries: bulk, a suspended membrane, a nanophononic crystal (NPC) with periodic internal cavities, and an NPM in the form of a nanopillared membrane (see geometric details in Fig. 1 and Appendix A). We assume the Stillinger-Webber interatomic potential in all models [38]. Moreover, the simulated nanostructures are assumed to exhibit atomically flat surfaces. This idealization is made to isolate the aforementioned mechanisms and exclude the effects of nonpristine phonon-boundary scattering. Future work may examine the effects of surface roughness consistent with experimental conditions.
@@ III Equilibrium configurational entropy
First, we study the available phase space by locally quantifying the configurational entropy at equilibrium. We use an approximated expression derived from an expansion of the configurational entropy in terms of the multibody correlation function [25]. In liquids, the two-body correlation term is enough to characterize most of the configurational entropy [31], and it has also been shown useful to study crystalline solids such as silicon [28, 44]. This term can be evaluated locally and for a given atom $i$ in the system reads
@@
In Fig. 2, we show local entropy maps for the different systems. Since we use the local density around the atom in Eq. (1), these calculations enable comparison of the available configurational phase space in different regions. We observe a larger accumulation of entropy close to the boundaries compared to the internal regions. Away from the boundaries, the crystalline order is enforced in all directions, thus strongly constraining each atom’s dynamics and reducing the size of the phase space of atomic configurations. In contrast, the atoms can accommodate a larger variety of microscopic configurations close to the boundaries due to the reduced number of neighbors. Consequently, the average equilibrium entropy $\bar{S}=\sum_{i}S^{i}/N$, where $N$ is the total number of atoms, increases with the surface-to-volume ratio, as clearly shown in the figure. This is consistent with the increased deviation of the dispersion from the bulk configuration. In addition, we observe that the entropy in the interior of the resonating nanopillars is significantly larger than in the interior of the membrane. We attribute this to the resonant modes in the nanopillar, which can correlate the dynamics of non-neighboring atoms, thus unlocking microstates that are not accessible in the membrane. Remarkably, this produces a significant increase in the size of the phase space in the NPM. We note that the observed differences in configurational entropy imply that each system needs to produce or exchange different amounts of entropy to reach equilibrium from given nonequilibrium conditions.
@@ IV Entropy production
A microscopic configuration of a crystal lattice can be interpreted in the reciprocal representation in terms of the phonon population, which only maximizes the entropy when accommodating to the Bose-Einstein distribution. Therefore, the out-of-equilibrium phonon evolution in the different systems can be investigated in terms of entropy evolution. Here we use steady-state nonequilibrium MD simulations, where the terminals of the system are connected to Langevin thermostats at $T_{\text{C}}=290$ and $T_{\text{H}}=310$ K, respectively. To attenuate unphysical coupling/interaction between thermostats, we consider five unit cells between the thermal baths (see Appendix D for details). In steady-state conditions, the entropy that the system produces internally is constantly exchanged with the thermostats, and quantifies the irreversible exchange of energy, or dissipation, to the system surroundings. According to the TUR [32], the thermodynamic uncertainty in terms of the energy current established in the system provides a lower bound on the entropy production $\Sigma_{\tau}$ during a time-window $\tau$:
@@
Figure 3: Nonequilibrium entropy production (a) bound, and normalized (b) mean and (c) variance of the energy transferred between thermostats $q_{\tau}$ through the different material systems over a time-window $\tau=$1 ns.
In such a stochastic interpretation of entropy production, the fraction of energy exchange with the thermostats that drive reversible phonon transitions in the system is not identified as energy dissipation and entropy exchange. Consistently, the entropy bound from the present MD simulations is smaller than the entropy exchange in the case of maximum energy dissipation (i.e., $\text{E}(q_{\tau})/T_{\text{C}}-\text{E}(q_{\tau})/T_{\text{H}}>\widehat{\Sigma}_{\tau}$) for all the cases considered, which in turn validates the TUR as a suitable tool in the present context [37, 48, 49]. Expression (3) is adequate to study the present systems far from equilibrium (where arbitrary non-Gaussian fluctuations may occur) beyond the linear-response interpretation [50]. Hence, $\widehat{\Sigma}_{\tau}$ is more robust than other metrics characterizing nanoscale nonequilibrium conditions such as the thermal conductivity, whose characterization requires strong assumptions. Nevertheless, it is illustrative to establish a relation between $\widehat{\Sigma}_{\tau}$ and an approximate bound for the thermal conductivity $k$ under the linear-response regime. Considering the entropy production rate in terms of the entropy fluxes $\Sigma_{\tau}/\tau=j/T_{\text{C}}-j/T_{\text{H}}$, with $j$ being the heat transfer rate, and assuming Fourier’s law, $\widehat{\Sigma}_{\tau}$ can be proportionally related to a lower bound on the conductivity [51],
@@
where $V$ is the system volume, $T$ is the average temperature, and $\nabla T$ is the thermal gradient imposed between the thermostats.
@@
In Fig. 3(a), we plot the entropy production bound $\widehat{\Sigma}_{\tau}/k_{\mathrm{B}}$ obtained from our nonequilibrium MD simulations considering $\tau=1$ ns. As demonstrated in Appendix D, this is the smallest sampling time for which we obtain $\text{Var}(q_{\tau})/\text{E}({q}_{\tau})^{2}\sim 1/\tau$. This relation is a consequence of the central limit theorem in the long sampling time limit [52], and allows us to characterize a bound for the entropy production rate $\widehat{\Sigma}_{\tau}/\tau$ independent of $\tau$. Since $q_{\tau}$ fluctuates both in time and between different simulation realizations, we perform the averages during 30 ns over five different simulations in each case. Analysis of the computational size effects and the influence of the temperature difference between thermostats is provided in Appendix D. As expected, the bulk is the system where entropy is produced most efficiently due to the absence of external constraints other than the thermostats. In contrast, the membrane exhibits a significant reduction in the entropy production rate. This is interpreted in terms of the local constraints to the phonon dynamics close to the free surfaces, which, in combination with the energy conservation restriction, inhibit some asymmetric transitions that are probable in the bulk case, thus reducing the overall probability of irreversible trajectories in the phase space that maximize the entropy. This is consistent with the lower $\widehat{\Sigma}_{\tau}$ observed in the NPC, where the addition of an internal cavity enhances the phonon-boundary interaction. In this case, however, we note that the entropy production reduction relative to bulk is too large to be solely attributed to more incoherent scattering due to an increase in the surface-to-volume ratio. Instead, the emergence of coherent interference of phonons, i.e., Bragg scattering, represents an additional constraint that can limit the efficient production of entropy. Finally, in the NPM, the addition of the nanopillars is shown to significantly reduce the production of entropy relative to the membrane even though no internal boundaries are included in the basis membrane, where the energy flows between the thermostats. Furthermore, the stronger the resonant effect due to a taller nanopillar, the lower the entropy the system produces. Considering the relation between $\widehat{\Sigma}$ and $\widehat{k}$, this finding suggests that, despite the strong anharmonicity governing a silicon crystal, introducing local resonances across the full spectrum is an efficient strategy to manipulate nanoscale thermal transport [11, 53].
@@
The rate of entropy production also provides a means to quantitatively distinguish between the roles of coherent dynamical effects versus noncoherent effects. In Figs. 3(b) and 3(c), we show $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$, respectively. Noticeably, boundary scattering in the membrane and the NPC reduces the entropy production mainly by reducing the average energy flux, which also implies a strong variance reduction, while, in the NPM, $\widehat{\Sigma}_{\tau}$ is reduced mainly due to an increase in flux fluctuations. Specifically, we observe that the flux variance increases monotonically by enhancing the resonant effect, even though the average energy current is not significantly modified. The strong current fluctuations established in the system are thus signatures of the prevalence of coherent phonon effects unlocked by resonances, while incoherent dynamics dominated by boundary scattering are characterized by a reduction in the energy transfer rate. According to Eqs. (2) and (3), these results imply that the energy transfer mechanisms in the presence of local resonances involve a reduced amount of irreversible internal transitions. This unique finding offers a broader view that nanoscale thermal energy conduction in general is not necessarily a fully irreversible process.

### v4

@@ Abstract
Thermal transport in crystals is influenced by chemistry, boundaries, and nanostructure. The anharmonic phonon band structure extracted from molecular-dynamics simulations provides an illuminating view of both the type and extent of prevalence of wavelike mechanisms underlying the transport, yet falls short of elucidating the nature of thermal evolution for different phonon regimes. Here we present an analysis framework for the characterization of the entropic signature of the mechanisms induced by boundaries and nanostructure, using both equilibrium and nonequilibrium atomistic simulations. Specifically, we examine the effects of phonon confinement, Bragg scattering, and local resonances on the configurational phase space in room-temperature nanostructured silicon, and quantify how each modifies the rate of entropy production and thermal relaxation. We reveal that the presence of phonon local resonances spanning the full spectrum enables a highly ordered regime of heat conduction to be approached, where irreversible evolution and entropy maximization are severely hindered by extensive mode hybridizations caused by the resonances. This unique regime of transport paves the way for ultra-precise phonon control for a wide range of applications in condensed matter physics.
@@ I Introduction
The second law of thermodynamics steers the spontaneous evolution of isolated material systems from arbitrary initial conditions towards the maximum entropy state corresponding to equilibrium. In crystal solids, Debye and Peierls, among others, conjectured that the anharmonic nature of atomic interactions combined with the presence of naturally-occurring lattice defects form the key causes behind the tendency of systems to thermally relax [1]. Consistent with this view, nonlinear phonon-phonon and phonon-defect scattering events are interpreted in the phonon gas picture to not necessarily conserve the lattice quasi-momentum, thus justifying on microscopic grounds the diffusion of vibrational energy along macroscopic thermal gradients. While all atomic interactions and motions are overseen by the Hamiltonian that governs a given system, the specific path and rate of evolution towards thermodynamic equilibrium may be affected by configurational constraints that drive certain mechanistic dynamical behavior. The inherent competition between mechanical response and the emergence of equilibrium were brought to light by pioneering computational experiments reported by Fermi and colleagues [2]. Lattice disorder and its impact on localization provides an example of a key mechanism that influences the evolution to equilibrium [3, 4]. More recently, a number of other mechanisms that may hinder thermodynamic behavior have been investigated, which include interaction with thermal reservoirs [5, 6], phonon confinement in low-dimensional systems [7, 8], and nonlinear interfacial effects [9].
@@
Nanostructuring offers yet another rich platform for the disruption of thermal evolution. When feature sizes are confined to the range of the phonon mean free path (MFP) at a given temperature, the intrinsic structural dynamics become a key factor in potentially influencing the path to equilibrium. In this context, a distinct mechanism that has been shown to cause transformative changes to the phonon band structure is atomic-scale resonance hybridizations [10]. The presence of sub-MFP resonating substructures generates localized phonon resonances that couple with the underlying heat-carrying phonons traveling in the host medium [11]. This phenomenon may in principle take place across the full phonon spectrum, thus bringing rise to a unique form of heat conduction that is based on resonant thermal transport [12]. In this regime, spectral energy density (SED) calculations obtained from equilibrium molecular-dynamics (MD) simulations have revealed distinctly appearing flat dispersion curves that couple with the underlying phonon dispersion throughout the full frequency range of atomic motion [13]. Nanostructures exhibiting this behavior, referred to as nanophononic metamaterials (NPMs), have recently been experimentally demonstrated [14]. Another wave-based phonon mechanism that also fundamentally influences the phonon band structure is Bragg scattering, which comes into effect with the presence of periodic features smaller than the phonon MFP [15, 16, 17, 18].
@@
The impact of the effects mentioned above on the emergence of thermal equilibrium has not yet been studied from a thermodynamics perspective. Yet, in other contexts, analysis techniques and experiments have been developed for the characterization of thermodynamic processes in terms of entropy evolution [19]--a perspective that in our view opens opportunities for exploring the interplay between structural dynamics and statistical mechanics [20, 21]. For example, generalized notions of entropy beyond local equilibrium have been applied towards the development of nanoscale thermal transport models not limited to diffusion [22, 23]. In addition, the equilibrium configurational entropy has been extracted directly from MD simulations [24, 25]. Noticeable applications include the characterization of phase transitions [26], defects [27, 28], biochemical phenomena [29], and the identification of universal scaling laws of transport properties in liquids [30]. Another key recent advancement is the emergence of thermodynamic uncertainty relations (TURs) [31, 32] built on fluctuation-dissipation theorems [33] as this provides new tools to relate the fluctuations of dissipative fluxes to the production of entropy in nonequilibrium processes [34, 35, 36].
@@
In this article, we quantify--for the first time--the entropic signature of non-diffusive heat flow in the presence of dynamical constraints created by material nanostructuring, and we do so using MD simulations. The resulting framework bridges the phonon dispersion properties of any given class of nanostructures with the thermodynamic picture and offers a new route to distinguishing between the coherent and noncoherent behavior of phonons. We consider a variety of crystalline nanoscale features that influence multiple phonon transport mechanisms at room temperature, including boundary scattering, geometric confinement, Bragg scattering, and local resonances. Considering silicon as an example constituent material, for each system we first obtain the phonon band structure by calculating the spectral energy density of atomic motion obtained by MD simulation where anharmonic effects are accounted for. The dispersion relations highlight the influence of the type of nanostructure on the wavelike nature of phonons. Next we characterize the configurational entropy distribution at equilibrium and uncover nonhomogeneous contributions influenced by the presence of the nanoscale features. We then combine nonequilibrium simulations and TUR to obtain a lower bound for the entropy production rate. This metric quantifies the impact of inherent dynamical phenomena on the production and exchange of entropy with the external environment, which allows us to characterize the strength of each configurational constraint in inhibiting the progression towards equilibrium. We find that NPMs are exceptional in this aspect owing to the unique ability of full-spectrum phonon-resonance hybridizations in hindering irreversible evolution. The existence of this kind of resonant phonon behavior, which manifests as a strongly fluctuating energy current, demonstrates that thermal energy currents do not necessarily correspond to entropy currents at the nanoscale. This brings to light a remarkable observation of a degree of reversibility in phonon transport in the presence of dynamical constraints.
@@
For purposes of comparison, we consider the following geometries: bulk, a suspended membrane, a nanophononic crystal (NPC) with periodic internal cavities, and an NPM in the form of a nanopillared membrane (see geometric details in Fig. 1 and Appendix A). We assume the Stillinger-Webber interatomic potential in all models [37]. Moreover, the simulated nanostructures are assumed to be defect-free and exhibit atomically flat surfaces. This idealization is made to isolate the aforementioned mechanisms and exclude the effects of impurity scattering and non-pristine phonon-boundary scattering (beyond the effects of atomic relaxation). Future work may examine the effects of material defects and surface roughness consistent with experimental conditions.
@@ III Equilibrium configurational entropy
First, we study the available phase space of atomic configurations by locally quantifying the configurational entropy at equilibrium. We use an approximate expression derived from an expansion of the configurational entropy in terms of the multibody correlation function [24]. In liquids, the two-body correlation term is enough to characterize most of the configurational entropy [30], and it has also been shown useful to study crystalline solids such as silicon [27, 43]. This term can be evaluated locally and for a given atom $i$ in the system reads
@@
In Fig. 2, we show local entropy maps for the different systems. Since we use the local density around the atom in Eq. (1), these calculations enable comparison of the available configurational phase space in different regions. We observe a larger accumulation of entropy close to the boundaries compared to the internal regions. Away from the boundaries, the crystalline order is enforced in all directions, thus strongly constraining each atom’s dynamics and reducing the size of the phase space of atomic configurations. In contrast, the atoms can accommodate a larger variety of microscopic configurations close to the boundaries due to the reduced number of neighbors. Consequently, the average equilibrium entropy $\bar{S}=\sum_{i}S^{i}/N$, where $N$ is the total number of atoms, increases with the surface-to-volume ratio, as clearly shown in the figure, for example, when comparing the membrane and NPC cases with the bulk case. This correlates with the increased deviation of the dispersion from the bulk configuration as observed in the corresponding phonon band structures in Fig. 1. In addition, we observe that the entropy in the interior of the resonating nanopillars is significantly larger than in the interior of the membrane. To quantify this difference, we calculated the average local configurational entropy in the center of the nanopillar (at a distance of 2 CC from its lateral boundaries) and compared it with the average local configurational entropy inside the membrane at a distance of 2 CC from the top and bottom boundaries. We found that the local average in the nanopillar is nearly 25% greater than that in the membrane. We attribute this to the resonant modes in the nanopillar, which can correlate the dynamics of non-neighboring atoms, thus unlocking microstates that are not accessible in the membrane. Remarkably, this produces a significant increase in the size of the phase space in the NPM. We note that the observed differences in configurational entropy imply that each system needs to produce or exchange different amounts of entropy to reach equilibrium from given nonequilibrium conditions.
@@ IV Entropy production
A microscopic configuration of a crystal lattice can be interpreted in the reciprocal representation in terms of the phonon population, which only maximizes the entropy when accommodating the Bose-Einstein distribution. Therefore, the out-of-equilibrium phonon evolution in the different systems can be investigated in terms of entropy evolution. Here we use steady-state nonequilibrium MD simulations, where the terminals of the system are connected to Langevin thermostats at $T_{\text{C}}=290$ and $T_{\text{H}}=310$ K, respectively. To attenuate unphysical coupling/interaction between thermostats, we consider five unit cells between the thermal baths (see Appendix D for details). In steady-state conditions, the entropy that the system produces internally is constantly exchanged with the thermostats, and quantifies the irreversible exchange of energy, or dissipation, to the system surroundings. According to the TUR [31], the thermodynamic uncertainty in terms of the energy current established in the system provides a lower bound on the entropy production $\Sigma_{\tau}$ during a time-window $\tau$:
@@
In such a stochastic interpretation of entropy production, the fraction of energy exchange with the thermostats that drive reversible phonon transitions in the system is not identified as energy dissipation and entropy exchange. Consistently, the entropy bound from the present MD simulations is smaller than the entropy exchange in the case of maximum energy dissipation (i.e., $\text{E}(q_{\tau})/T_{\text{C}}-\text{E}(q_{\tau})/T_{\text{H}}>\widehat{\Sigma}_{\tau}$) for all the cases considered, which in turn validates the TUR as a suitable tool in the present context [36, 47, 48]. Expression (3) is adequate to study the present systems far from equilibrium (where arbitrary non-Gaussian fluctuations may occur) beyond the linear-response interpretation [49]. Hence, $\widehat{\Sigma}_{\tau}$ is more fundamental, and robust, than other metrics characterizing nanoscale nonequilibrium conditions such as the thermal conductivity $k$, which is generally a phenomenological concept. Quantification of $k$, using the Green-Kubo method [50, 51] or directly by nonequilibrium simulations [52, 53], for example, requires the assumption of Fourier’s law--which is not necessarily valid at the nanoscale [54]. In contrast, quantification of $\widehat{\Sigma}_{\tau}$ does not require imposing any constitutive relation between the heat flux and the thermal gradient, which extends the validity of the analysis to complex nanoscale conditions well beyond the diffusive transport regime. Nevertheless, it is illustrative to establish a relation between $\widehat{\Sigma}_{\tau}$ and an approximate bound for the thermal conductivity $k$ for cases displaying a linear response. If we consider the entropy production rate in terms of the entropy fluxes $\Sigma_{\tau}/\tau=j/T_{\text{C}}-j/T_{\text{H}}$, with $j$ being the energy dissipation rate, and assume Fourier’s law, $\widehat{\Sigma}_{\tau}$ can be proportionally related to a lower bound on the thermal conductivity [55],
@@
where $V$ is the system volume, $T$ is the average temperature, and $\nabla T$ is the thermal gradient imposed between the thermostats [^1] For the bulk case, the thermal conductivity lower bound from Eq. (4) is approximately five times smaller than what is predicted by comparing the heat flux with the thermal gradient in nonequilibrium simulations..
Figure 3: Nonequilibrium entropy production for four material systems considered with computational models displayed in (a). Shown are the (b) bound, and normalized (c) mean and (d) variance of the energy transferred between thermostats $q_{\tau}$ through each material system over a time-window $\tau=$1 ns. In the NPM case, a decrease in the entropy production bound is observed when the nanopillar height is increased--a behavior associated with the energy current fluctuations induced by the resonant dynamics of the nanopillars.
@@
In Fig. 3(b), we plot the entropy production bound $\widehat{\Sigma}_{\tau}/k_{\mathrm{B}}$ obtained from our nonequilibrium MD simulations considering $\tau=1$ ns. As demonstrated in Appendix D, this is the smallest sampling time for which we obtain $\text{Var}(q_{\tau})/\text{E}({q}_{\tau})^{2}\sim 1/\tau$. This relation is a consequence of the central limit theorem in the long sampling time limit [57], and allows us to characterize a bound for the entropy production rate $\widehat{\Sigma}_{\tau}/\tau$ independent of $\tau$. Since $q_{\tau}$ fluctuates both in time and between different simulation realizations, we perform the averages during 30 ns over five different simulations in each case. Analysis of the computational size effects and the influence of the temperature difference between thermostats is provided in Appendix D. As expected, the bulk is the system where entropy is produced most efficiently due to the absence of external constraints other than the thermostats. In contrast, the membrane exhibits a significant reduction in the entropy production rate. This is interpreted in terms of the local constraints to the phonon dynamics close to the free surfaces, which, in combination with the energy conservation restriction, inhibit some asymmetric transitions that are probable in the bulk case, thus reducing the overall probability of irreversible trajectories in the phase space that maximize the entropy. This is consistent with the lower $\widehat{\Sigma}_{\tau}$ observed in the NPC, where the addition of an internal cavity enhances the phonon-boundary interaction. In this case, however, we note that the entropy production reduction relative to bulk is too large to be solely attributed to more incoherent scattering due to an increase in the surface-to-volume ratio. Instead, the emergence of coherent interference of phonons, i.e., Bragg scattering, represents an additional constraint that can limit the efficient production of entropy. Finally, in the NPM, the addition of the nanopillars is shown to significantly reduce the production of entropy relative to the membrane even though no internal boundaries are included in the basis membrane, where the energy flows between the thermostats. Furthermore, the stronger the resonant effect due to a taller nanopillar, the lower the entropy the system produces. Considering the approximate relation between $\widehat{\Sigma}$ and $\widehat{k}$, this finding suggests that, despite the strong anharmonicity that governs a silicon crystal, introducing local resonances across the full spectrum is an efficient strategy to manipulate nanoscale thermal transport [10, 58].
@@
The rate of entropy production also provides a means to quantitatively distinguish between the roles of coherent dynamical effects associated with any existing wave behavior of phonon motion versus noncoherent effects associated with nonlinear phonon interactions and scattering from non-pristine surfaces. In Figs. 3(c) and 3(d), we show $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$, respectively. Noticeably, boundary scattering in the membrane and the NPC reduces the entropy production mainly by reducing the average energy flux, which also implies a strong variance reduction, while, in the NPM, $\widehat{\Sigma}_{\tau}$ is reduced mainly due to an increase in flux fluctuations. The strong energy current fluctuations established in the NPM models are thus signatures of the prevalence of coherent phonon effects generated by the nanopillar resonances, while incoherent dynamics dominated by anharmonic and boundary scattering are characterized solely by a reduction in the energy transfer rate. Precisely, in the NPMs we observe that the flux variance increases monotonically by enhancing the resonant effect, even though the average energy current is not significantly modified. According to Eqs. (2) and (3), this implies that the energy transfer mechanisms in the presence of local resonances involve a reduced amount of irreversible internal phonon transitions. This unique finding offers a broader view that nanoscale structures accommodating strong dynamical constraints can induce some degree of reversible evolution while sustaining thermal energy currents.